How do you get the answer to x squared minus 14<15
Answer by stanbon(48569)

How do you get the answer to x squared minus 14<15
-------------------
x^2 - 14 < 15
x^2 < 29
-------------
1st: Solve the equality x^2=29
x = sqrt(29) or x = -sqrt(29)
----------------------------------
2nd: Draw a number line and plot -sqrt(29) and +sqrt(29)
----------------------------------
3rd: Test x values from each of the three number-line intervals
to see where the solutions for the inequality lie.
Check x = -10: (-10)^2 < 29 ; false
Check x = 0 : (0)^2 < 29 ; true
Check x = +10: (10)^2 < 29 ; false
------------------------------------------
Final Solution:
-sqrt(29) < x < +sqrt(29)
====================================
Cheers,
Stan H.